every evening jenna empties out her pockets and puts her change in a jar. at the end of the week she counts her money. one week she had 40 coins, all of them dimes and quaters. when she added them up she had a total of &7.75.
how many dimes did jenna have??
Answer #1
Let Q = the number of quarters and D = the number of dimes. Then: 1) Q+D = 40 "One week she had 40 coins, all of them dimes and quarters...." 2) 0.25Q+0.1D = $7.75 "...she had a total of $7.75." Rewrite equation 1) as: D = 40-Q and substitute into equation 2). 2a) 0.25Q+0.1(40-Q) = 7.75 Simplify the left side. 2b) 0.25Q+4-0.1Q = $7.75 Subtract 4 from both sides. 2c) 0.25Q-0.1Q = 3.75 Combine the like-terms on the left side. 2d) 0.15Q = 7.75 Divide both sides by 0.15 2e) Q = 25 and: D = 40-Q D = 40-25 D = 15 Jenna had 25 quarters and 15 dimes.
answered by: delmy